Problem Solving Set 20

23 July 2012

1. A, B are square complex matrices and rank(AB −BA) = 1. Show that (AB − BA)^{2} = 0.

2. Consider the equation

f′(x) = f(x + 1).

Is there a solution with f(x) → ∞ as x → ∞?